Setup on argos source with LBTI: Baysides: 63.3, 25.76, 73.8 M2 X2.65 Y1.9 Z1.03 RX191.85 RY-171 RZ0 M3 Absolute -20 -20 0 2.4

1:31: Tried to close loop but failed and put adsec in failure statue during center pupils
AdSecArbitrator _74946|INF|      1012|2021-01-22 01:20:02.959507|             MAIN > Received cmd: LoadRecMatTag ("KL_v20", "LBTI_IIR_b1_500m1000", "LBTI_IIR_b1_500m1000", 0)
AdSecArbitrator _74946|INF|      1013|2021-01-22 01:20:02.959583|   COMMANDHANDLER > FSM (status: AOSet) has received command 2022 (SetRecMat)
AdSecArbitrator _74946|ERR|      1014|2021-01-22 01:20:04.243580|              IDL > Command returns error : Timeout error
AdSecArbitrator _74946|INF|      1015|2021-01-22 01:20:04.243766|   COMMANDHANDLER > Command execution failed: Fsm forced to Failure
Fri Jan 22 01:24:23.494 2021 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[1382] - [AOException] SelectFocalStation: WARNING - adsecarb:Illegal command for state Failure (Set focal station)
Fri Jan 22 01:25:07.126 2021 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[1384] - [AOException] SelectFocalStation: WARNING - adsecarb:Illegal command for state Failure (Set focal station)
...
Fri Jan 22 01:29:51.130 2021 aos.info - FLAO: Adsec 672b-left recovery from FAILURE started.
Recovered then instead of running preset through the AOSGUI I manually closed no issue. I'm not sure what triggered the inability to load the Rec matrix tag. I initially thought that maybe there was an issue with the Focal station not being set (disproven later tonight), so set the focal station manually to Center.

Preset through AOSGUI for 10.5 mag no issue.

Playing with NCPA to try and tweak the PSF on the LMIRCam images, although I'm not quite in peak form today for this. I caused a failure by applying too high forces to the mirror with too much NCPA.
Fri Jan 22 02:07:39.037 2021 aos.info - LBTI WFS daemon went off
Fri Jan 22 02:07:40.579 2021 aos.info - TTF offload command end
Fri Jan 22 02:07:45.969 2021 aos.info - FLAO Alert [ code:0] Error
Fri Jan 22 02:07:46.318 2021 aos.info - Updated variable adamhousekeeper.L.COIL_STATUS: 0
...
Fri Jan 22 02:09:31.453 2021 aos.info - FLAO: Adsec 672b-left recovery from FAILURE started.
Recovered and just after.... 2:20 Following error on SX hexapod left leg. Steve is recovering IT#5935:Running closed dome AO testing with LBTI, closed loop:
Fri Jan 22 02:17:53.188 2021 left Secondary moveAbs failed because Mirror not in position
Fri Jan 22 02:17:57.628 2021 left Secondary moveAbs failed because Mirror not in position
Leg#4 following error. 02:22UT Opened the loop, and Steve ran these commands in OSSGUI > SXM2 subGUI, as before: #4j/ #4k #4j/ Open brakes

2:48 GOpt gui is showing stale NCPA values after all of the various troubleshooting efforts. I closed it but cant get it to come back up. Searched for running process but can't find any lingering.

The numarray module will be dropped in Numpy 1.9
Traceback (most recent call last):
  File "/home/sxwunit/aoroot/bin/goptgui.py", line 356, in <module>
    main = gopt(qApp)
  File "/home/sxwunit/aoroot/bin/goptgui.py", line 145, in __init__
    self.getz0()
  File "/home/sxwunit/aoroot/bin/goptgui.py", line 347, in getz0
    self.z0 = self.asm.get_z0()
  File "/home/sxwunit/aoroot/lib/python/AdOpt/asm.py", line 49, in get_z0
    self.ex( cmd, timeout=5000)
  File "/home/sxwunit/aoroot/lib/python/AdOpt/asm.py", line 35, in ex
    return self._app.IDLCmd( cmd, self._server, timeout)
  File "/home/sxwunit/aoroot/lib/python/AdOpt/thAOApp.py", line 643, in IDLCmd
    answer = self.SendMsg(idlctrl, AOMsgCode ['IDLCMD'], cmd, timeout=timeout, want_answer = True)
  File "/home/sxwunit/aoroot/lib/python/AdOpt/thAOApp.py", line 340, in SendMsg
    return thrdlib.thWaitMsg( seqn = seqn, timeout = timeout)
thrdlib.error: thWaitMsg error(-5001): Timeout error
Segmentation fault (core dumped)
I verified no other goptgui open or process running…

w-stop_all and w_start_all and gopt is happy again. But looks like Steve is happy with no NCPA after all.

Closed loop with Bin2. Optimize gain parameters are nowhere near set for bin 2, (and probably not for bin3). Should we steal some from LUCI soul? Closed and optimized by eye easily enough.

Set the focal station to null and sent a preset AO from the AOSGUI. Verified the PresetAO does set the focal station as expected. Not sure what happened earlier.

Shutting down, parking the arsgos source. Shutter fully closed and off. disabled rr mode. Now we wait for humidity to drop.

-- JenniferPower - 22 Jan 2021
Topic revision: r3 - 25 Jan 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback